OpenStudy (anonymous):

The sum of two consecutive intergers is 153. Which equation can be used to find the first interger n?

hartnn (hartnn):

Let the smaller number be 'n' consecutive means next number = n+1 so, sum = n+(n+1) now can you find the equation ?

OpenStudy (anonymous):

@hartnn is right, but i would use common sense

hartnn (hartnn):

like dividing 153 by 2 ??

OpenStudy (anonymous):

i.e. guess and check try 73 + 74, no too low

hartnn (hartnn):

153/2=76.5 so, 76+77

OpenStudy (anonymous):

math teachers like to use equations, but don't forget you are a thinking human being with a brain. if someone asked you outside of math class to come up with two numbers whose sum was 153 and that were one apart, i am going to bet that within one minute you could come up with 76 + 77 = 153

OpenStudy (anonymous):

let the number be x the other consecutive is x+1 so x+x+1=153 2x=152(subtract both side 1) x=76(divide both side by 2) so the numbers are 76 and 77 That's it!!!!!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait it has answer choices

OpenStudy (anonymous):

A. 2n + 1 = 153 C. n + 1 =153 B. 2n = 153 D. 2n + 2 =153

OpenStudy (anonymous):

can anyone help me yes no maybe

hartnn (hartnn):

n+(n+1) = 2n+1 =153